Tax operating loss

Fore Farms reported a pretax operating loss of $180 million for financial reporting purposes in 2021. Contributing to the loss were (a) a penalty of $8 million assessed by the Environmental Protection Agency for violation of a federal law and paid in 2021 and (b) an estimated loss of $10 million from accruing a loss contingency. The loss will be tax deductible when paid in 2022.
 
The enacted tax rate is 25%. There were no temporary differences at the beginning of the year and none originating in 2021 other than those described above.

 
Required:
1. Prepare the journal entry to recognize the income tax benefit of the net operating loss in 2021.
2. What is the net operating loss reported in 2021 income statement?
3. Prepare the journal entry to record income taxes in 2022 assuming pretax accounting income is $205 million. No additional temporary differences originate in 2022.
Answer #1

Req.1 Journal Entry:

DateAccount NameDebitCredit
1.Deferred tax asset$43 m

To Income tax expense
$ 43 m

( to recognize income tax benefit)

( $ 180mn - 8mn = $ 172mn

$ 172 mn * 25%)

Req. 2

Net Operating Loss$ 137 mn

(- 180 mn + 43mn)

Req. 3

DateAccount NameDebitCredit
2022Income Tax Expense$51.25

To Deferred Asset
$41.50

To Income Tax Payable
$9.75

Working:

1) $205 - $10= $ 195mn

2) $ 195 *20% = $39 ; $ 195 *80% = $ 156

3) $ 39 ×25% = $ 9.75

4) 172 - ( 156 + $ 10)

= 6

5) $6mn* 25%= $ 1.5 mn

6) $ 43mn - $ 1.5 mn = $ 41.5


answered by: Astars